What is the concentration of 25.00 mL of an unknown monoprotic acid if 18.24 mL of standardized 0.125 M sodium hydroxide solution was required to reach the equivalence point of the titration?
Answer #1

Monoprotic acid = HA

Sodium hydroxide = NaOH

HA + NaOH -----> NaA + H2O

M1V1 = M2V2

M1 = ?, V1 = 25.00 mL

M2 = 0.125M, V2 = 18.24 mL

M1*25.00 = 0.125*18.24

M1 = (0.125*18.24)/25.00

     = 0.0912 M